INAF queue observing, 2015 Nov 29-30 
Observer: F. Cusano, A. Rossi 
Telescope Operator:  Steve Allanson 
Support Astronomer: Dave Thompson (on call) 
Instrument:  LUCI/MODS/LBC
  Summary 
At the sunset the sky is clear. Wind at the limit (18m/s) 
Bad seeing at the begining of the night up to 2.5", decreased 
during the night (7UT) up to 1". Problem with the pointing of the telescope 
solved in ~1h.
Night duration (12\xB0 el twilight) 18:20-6:00 ( 02:30-11:50 UT) = 
11 h 40m
Observing time: 
10 hrs 
40 m* (81%)
Weather time loss:   
(0%)
technical time loss  
1h (9%)
  Data Summary 
	
		
			|  Project_RUN  | 
			 Objects  | 
			 Filter  | 
			 ExpTime  | 
			 Notes  | 
		
	
	
		
			|  Dark LUCI  | 
			 Ragazzoni, Guaita  | 
			    | 
			    | 
			    | 
		
		
			|  Rossi Too  | 
			 GRB151027A  | 
			 LUCI J, LBC r  | 
			 40m  | 
			 parallel LUCI2+LBC-B  | 
		
		
			|  Cusano  | 
			 AndXXI, AndXIX, field  | 
			 B,V  | 
			 30m  | 
			 AndXXI _ob AndXIX _ob 3 and 4, field_2  | 
		
		
			|  Dalessandro  | 
			 N188  | 
			 B,V, I  | 
			 1h  | 
			 completed  | 
		
		
			|  Guaita  | 
			 M0717  | 
			 H2, Br_Gam  | 
			 1h, 2h  | 
			 some images out of constraint for seeing  | 
		
	
  Details 
UT date 20151129 
22:38 Dark LUCI:  
luci2.20151130.0002-0022 Guaita Dark,  
luci2.20151130.0023-0033 Ragazzoni dark 
    
luci2.20151130.0034-0054 Rossi dark
01:37 starting parallel observation with LUCI2 and LBC-B, Target of opportunity  Rossi 
-  imaging luci2.20151130.0055-0095==
  -  dofpia
  -  LBC-B lbcb.20151130.014618.fits jump in the image, not good; the other are fine
 
 
02:13 swith to MODS
02:22 slewing to standard feige110, we can not perform the observations of the standard because is facing the wind
02:27 problem with the pointing; telescope problem with the balancing
3:38 problem solved by Kevin, we slew to Saracco field 
-  acqMODS2 mods.4.XLSms2.acq 
mods2r.20151130.0002-0006
  -  seeing 2", out of the spec we give up
 
 
4:04 switch to LBC
4:18 slewing to Cusano program 
-  dofpia
  -  dimm seeing 1.7"
  -  copointing
 
 
4:34 executing  
AndXXI _3.ob 
-  dimm seeing 2.2"
  -  measured seeing on the image 1.8"
 
 
4:44 executing 
AndXIX _3.ob
  * seeing 1.6 on the images
4:53 executing And_field_2.ob 
 
5:13  excuting 
AndXXI _4.ob 
 
5:21 executing 
AndXIX _4.ob 
-  dimm 1.5"
  -  seeing 1.4" on the images
 
 
5:30 slewing to Dalessandro program 
 
5:43 executing N188_USNO_P5.ob
5:45 executing N188_BVI_1.ob 
-  dimm 1.5"
  -  seeing on the images 1.4"
 
 
5:54 executing N188_BVI_1b.ob
6:01 executing N188_BVI_2.ob 
-  seeing 1.3" red channel, 1.5" blue channel measured on the images
 
 
6:11 executing N188_BVI_2b.ob
6:18 dofpia
6:24 executing N188_BVI_3.ob 
-  dimm 1" ! also on the images
  -  stars elongated in the blue, we run again dofpia
 
 
6:36 executing N188_BVI_3b.ob
6:42 running copointing again
6:47 executing N188_BVI_4.ob
6:56 executing N188_BVI_4b.ob
7:02 executing N188_BVI_5.ob
7:12 executing N188_BVI_5b.ob
7:20 switch to LUCI
7:48 imaging program of Guaita, Guaita Br_gam 
luci2.20151130.0115-0175 
-  at the start seeing 1.4" (on the GCS)
  -  seeing improving 1" at 8:01
  -  seeing again not so good 1.4" at 8:20
  -  seeing 1" at 8:38, the images look good
  -  performing some test with LBC-B
  -  seeing measured on the image 0.9" at 9:03
 
 
9:30 Guaita H2, 
luci2.20151130.0176-0235 
-  seeing 0.8" measured on the images
  -  10:39* seeing 1 measured on the images
  -  11:13 seeing 0.9 on the images
 
 
11:20 Guaita Br_Gam, 
luci2.20151130.0236-0295 
-  seeing 0.8" on the images
 
 
13:00 starting MODs calibration of Saracco
    * slitless flat red grating 
mods2r.20151130.0007-00011
    * 5 arcsec flat stopped problem with the grabber of MODS2